# AudioMoG: Guiding Audio Generation with Mixture-of-Guidance

### v1

@@ Appendix I More generated samples
We shown more text-to-audio and video-to-audio generation results in Figure 8 and 9, respectively. As shown in Figure 8, HG and PG consistently achieve more accurate harmonic modeling and superior temporal alignment compared to CFG in the first example. Their outputs exhibit well-defined harmonic stacks and consistent overtone structures, even in complex or polyphonic cases in the third example. Moreover, HG and PG maintain precise timing across events, effectively capturing the onset and duration of audio elements in the second example. In contrast, CFG often fails to organize harmonics coherently and produces temporally smeared results. These comparisons clearly illustrate the advantage of our methods in reinforcing both spectral clarity and temporal fidelity. In Figure 9, HG generates significantly clearer high-frequency content and overall higher-quality audio compared to CFG in the first example. The resulting audio exhibits more natural brilliance and detail in the upper frequency range, enhancing perceptual realism. For the second example, HG demonstrates superior temporal alignment, accurately synchronizing audio events with visual cues, while CFG shows noticeable temporal drift and inconsistent timing. These comparisons further highlight the strengths of our method in improving both spectral resolution and temporal coherence in V2A generation. For more generated samples, please refer to our demo page: https://audiomog.github.io.
@@
  • EDM2 models (Karras et al., 2024b; a): Creative Commons BY-NC-SA 4.0 license
  • Stable Audio Tools (Evans et al., 2024): MIT license
  • AudioLDM-Eval (Liu et al., 2023): MIT license
  • CLIP (Radford et al., 2021): MIT license
  • Diff-foley models (Luo et al., 2023): Apache-2.0 license
  • VTA-LDM models (Xu et al., 2024): Apache-2.0 license
  • FoleyCrafter models (Zhang et al., 2024): Apache-2.0 license
  • V2A-Mapper models (Wang et al., 2024a): Creative Commons BY-NC-ND 4.0 license

### v2

@@ Appendix I More generated samples
We shown more text-to-audio and video-to-audio generation results in Fig. 7 and 8, respectively. As shown in Fig. 7, HG and PG consistently achieve more accurate harmonic modeling and superior temporal alignment compared to CFG in the first example. Their outputs exhibit well-defined harmonic stacks and consistent overtone structures, even in complex or polyphonic cases in the third example. Moreover, HG and PG maintain precise timing across events, effectively capturing the onset and duration of audio elements in the second example. In contrast, CFG often fails to organize harmonics coherently and produces temporally smeared results. These comparisons clearly illustrate the advantage of our methods in reinforcing both spectral clarity and temporal fidelity. In Fig. 8, HG generates significantly clearer high-frequency content and overall higher-quality audio compared to CFG in the first example. The resulting audio exhibits more natural brilliance and detail in the upper frequency range, enhancing perceptual realism. For the second example, HG demonstrates superior temporal alignment, accurately synchronizing audio events with visual cues, while CFG shows noticeable temporal drift and inconsistent timing. These comparisons further highlight the strengths of our method in improving both spectral resolution and temporal coherence in V2A generation. For more generated samples, please refer to our demo page: audiomog.github.io.
@@
  • EDM2 models [30, 31]: Creative Commons BY-NC-SA 4.0 license
  • Stable Audio Tools [13]: MIT license
  • AudioLDM-Eval [42]: MIT license
  • Diff-foley models [46]: Apache-2.0 license
  • VTA-LDM models [59]: Apache-2.0 license
  • FoleyCrafter models [61]: Apache-2.0 license
  • V2A-Mapper models [55]: Creative Commons BY-NC-ND 4.0 license